### v1

@@ 4 Results
The analysis and synthesis of the data extracted from our finalized 63 primary studies are reported in the following subsections. The results discussed in this section are derived by synthesizing the data with minimal interpretations directly extracted from the selected primary study set.
@@ 4.1 Demographic Data
Of the 63 primary studies selected for this SLR, 36 (57.1%) were conference papers, 25 (39.6%) were journal papers, and 2 (3.1%) were workshop papers. Figure 3 shows the yearly distribution and publication type distribution of the selected papers. The selected primary studies were published between 2011 and 2023 era even though we did not limit our search to a specific period. None of the included studies were published in 2012. There is an overall increase in the number of publications from 2011 to 2018. Then, in 2019 and 2020, there was a drop in the number of publications. The COVID-19 outbreak might have caused the sudden publication drop seen in 2020. However, it is not certain. Overall, 50% of the selected primary studies set were published from 2021 to 2023.
@@
Contributor’s social characteristics S8, S12, S34, S36, S38, S53
Previous Contributions Skills, Expertise S2, S5, S10, S25, S27, S28, S29, S32, S37, S39, S40, S41, S42, S43, S45, S46, S52, S55
@@
Historical behavior sequences S5, S7, S8, S10, S17, S18, S19, S20, S21, S22, S23, S24, S32, S36, S47, S50, S52, S55, S60
@@
    • Optimization - Genetic Algorithm - Genetic Algorithms are based on evolutionary principles such as natural selection and survival of the fittest [24]. Study S14 transforms the issue of recommending tasks to new developers into a multi-optimization problem. Then, it employs a topic-sampling based genetic algorithm to recommend tasks.
@@
Increasing performance, accuracy, and optimizing solutions: The performance, accuracy, and system optimizations in the selected primary studies were mainly achieved by using two or more characteristics mentioned in Section 4.2.1. For example, studies S8, S12, S34, S36, S38, and S53 used software practitioner’s social connections with fellow software practitioners along with other characteristics to improve the accuracy of the systems. According to Study S12, the use of social metrics in their recommendation system improved the precision of the used models up to 15.82%. Further, some studies (e.g., S5, S7, S27) used hybrid recommendation systems (mentioned in section 4.2.2) to mitigate the limitations that occurred when using one recommendation approach.
@@
Study S28 solves the “cold start” problem by modeling worker’s preferences over multiple dimensions. Study S44 recommends tasks to newcomers by integrating repository history data (such as the number of prior commits made by the issue reporter at the time of the issue’s creation) and global GitHub profiles for all issue reporters and repository owners (e.g., stars received, commits). Study S59 recommends tasks to cold-start users by analyzing possible features and the characteristics of the newcomers’ chosen first issues. A new cross-platform recommendation approach for software crowdsourcing was introduced in S61 by transferring data and knowledge from other popular software crowdsourcing platforms. Study S61 adopts a transfer learning approach to solve the platform “cold start” problem.
@@
    Limitations in the approaches (7 papers): One of the key limitations identified in the proposed approaches is the lack of generalizability of the solution. The recommendation model in Study S3 primarily follows a content-based approach, limiting its ability to suggest issues to developers based solely on past contributions. It lacks the capability to provide more insightful recommendations that consider serendipity aspects across various components of the same product. Hence, the recommendations could be narrow. The NLP suggestions used in Study S12 might not work as expected in other languages like C++. Therefore, the approach cannot be utilized in projects using different programming languages. Whereas in Study S9, the projects within the scope of the study exhibit diverse practices, backgrounds, resources, and goals. These factors are not adequately considered during the model-building process. Hence, the recommendation system proposed in the study may not be applicable or valid across a broader range of projects with different contexts. Another key limitation is that some systems cannot recommend tasks to newcomers ( “cold start” problem). In Studies S2, S14, S15, and S17, the proposed recommendation systems cannot recommend tasks to newcomers. As mentioned earlier in section 3.1, human factors play a major role in the software engineering context. However, only 24 primary studies have considered human factors in their recommendation systems. These human factors are analyzed in section 4.4. 39 primary studies ignored human factors in their recommendation systems. Some primary studies that ignored human factors are, S3, S6, S7, S10, S11, S14, S17, S20, 21, S22, etc.
@@ 4.3 RQ2:What are the common types of tasks that are recommended to software practitioners?
4.4 RQ3: To what extent are human factors used to recommend tasks in software repositories?
To answer this research question, we analyzed the human factors (i.e., D09 and D10 in LABEL:table:7) currently used in the existing recommendation systems and the human factors suggested to be included as future improvements to the recommendation systems. Table 10 shows these human factors.
@@ 4.4 RQ3: To what extent are human factors used to recommend tasks in software repositories?
Studies S4, S57, and S58 investigated the impact of personality on task selection in crowdsourcing software development. In study S4, an empirical experiment was carried out to assess the impact of personality on task selection based on the three key components of a task: type, money, and time [21]. Study S57 was carried out to develop a task assignment algorithm for crowdsourcing software development that takes task types and personality types (e.g., Introversion and Extroversion, etc. ) into account. The proposed method functions in the same way as the open call format, but it incorporates personality-based categorization [22]. To identify the personality types of software practitioners, the Myers-Briggs Type Indicator (MBTI) was used [26]. MBTI is a popular tool used to analyze personality types in the workplace which measures a person across four dimensions: energizing, attending, deciding, and living [27]. Study S58 was carried out to find the relationship between testing technique, tester, and personality type and identify which personality types of testers are more competent at a particular testing approach [28]. They also used the MBTI to identify the personality types [27].
@@
Demographic information, such as age, location, education, etc., are currently included in Studies S15 and S28, while Study S38 suggested considering the geographic region in the future. Study S15 modeled developer and project profiles by analyzing their basic information and mining user online behavior logs. Then they filtered out the projects based on the location, age, and education level requirements. This filtration process was done to speed up the recommendation process. It has been observed that workers from certain countries tend to prefer collaborating with task posters from specific other countries. For example, workers from India often prefer to work with task posters from the US and the UK [29]. Hence, Study S28 captured a worker’s preference for collaborating with task posters from certain countries. They extracted the worker’s location, education, etc. from the worker’s profile information.
@@
Skill preference S1, S5, S15, S16, S32, S59 N/A
Social and behavioral preferences S8, S10, S43 S5
@@ 4.5 RQ4:What platforms are currently being used to recommend crowdsourced software engineering tasks?
GitHub is the most popular platform used in the primary studies. It is used in 36 out of 63 studies (57.1% of the studies). It is a developer-based platform to build, scale, and deliver secure software [31]. Some of the Github projects used in the existing studies are JabRef, PowerToys, and Audacity. A further analysis of these projects is available in table LABEL:github_project_scales.
@@
RQ4 Summary: We observed that GitHub is the most popular (57.1%) crowdsourced software engineering platform among the selected primary studies. Further, we observed that the majority (55%) of the GitHub projects that were used to recommend tasks were large-scale projects with more than 500 contributors. Further, Bugzilla, Topcoder, and Utest are some other popular CSE platforms (see Table LABEL:platforms_list).
@@ 4.6 RQ5:Which features of platforms can be used to recommend crowdsourced software engineering tasks?
Table 13 shows the features and the different attributes of the features of crowdsourcing platforms used to extract data.
@@
User activities such as starring, forking, watching, bidding, were extracted in 22 primary studies. This is the second most popular feature used in primary studies. For example, Study S7 extracted starring, and forking repositories in the GitHub platform to identify user preferences.
Pull requests are the third most popular feature used to extract different attributes. It was used in 17 primary studies. Commit message text, names of the files changed in the pull request, and pull request comments were extracted from the pull requests. For example, Study S1 extracted the name of the files changed, and commit messages from the pull requests to identify the relevant domain for the issues.
Comments were used in 12 primary studies. commenter’s name and the text were extracted from the comments. Study S12 extracted commenter names to get the total number of commenters which they used as a social metric in their recommendation system.
@@
RQ5 Summary: Issue or task description is the most popular feature used to extract data in primary studies. Further, pull requests, comments, labels, tags, etc. are other popular features used to extract data (see Table 13) From those features, different attributes such as commit message, task type, APIs used, etc were extracted.
@@
User activities Fork, star, watch, bid, etc S5, S7, S16, S18, S19, S20, S21, S22, S23, S24, S32, S34, S35, S36, S44, S47, S50, S51, S56, S57, S62, S63
Pull Requests Commit messages S1, S10, S11, S12, S18, S20, S33, S36, S41, S44, S53, S59
@@ 5 Threats to Validity
Search String. To address the common threat of overlooking relevant papers in any systematic literature review (SLR), as detailed in Section 3.4, we employed six prominent digital libraries for paper retrieval. We contend that utilizing Scopus, recognized as the largest indexing system with the most extensive search capabilities among other digital libraries, allowed us to enhance the inclusiveness of relevant studies. Furthermore, we implemented two other strategies to mitigate potential search-related threats. We refined the search string iteratively based on pilot searches, ensuring careful testing before executing the search for relevant papers. We also employed a snowballing technique that involves manual searches of references in selected papers during the second phase of the search process, aiming to identify additional related papers.
Study Selection. The process can be susceptible to researchers’ subjective judgment regarding whether a paper meets the selection criteria for inclusion or exclusion. To mitigate potential biases in study selection, we adhered rigorously to the predefined review protocol. We recorded reasons for inclusion and exclusion to facilitate ongoing internal discussions between the first and second authors regarding any papers that raised doubts about their inclusion or exclusion decisions. Initially, the first two authors validated the inclusion and exclusion criteria on a small subset of primary studies, resolving any disagreements through discussion.
Data Extraction. Bias from researchers during data extraction poses a basic threat to any SLR, potentially impacting its results. To address this concern, we implemented several measures. We first developed a data extraction form (see Table 7) to ensure consistent extraction and analysis of data, aligned with the research questions of the SLR. Given that the first author conducted a significant portion of the data extraction, regular discussions were organized with the second author to address any uncertainties and rectify discrepancies in the extracted data. Finally, a subset of the extracted data underwent verification by the second and third authors.
Data Synthesis. We made an effort to minimize our subjective judgments and interpretations when analyzing the gathered information. All authors of this SLR discussed how to effectively categorize and present data in multiple rounds of discussion to minimize possible biases during data analysis and synthesis.
@@ 6 Discussion and Future Research Directions
    Software is developed by humans, read by humans, and maintained by humans [38]. Human factors play a crucial role in software development [39]. Human factors affect software development team members, their activities, and ultimately the quality of the product [15]. Human factors have been integrated into some existing primary studies, and several studies plan to incorporate them as future enhancements to their recommendation systems. We analyzed the human factors currently being used or suggested to be used in the existing systems in Table 10. For example, Studies S9 and S59 used software practitioners’ sentiment data in their recommendation system. Study S29 considers software practitioners’ motivation when recommending tasks, whereas Studies S18 and S28 plan to consider motivation as a future improvement. Studies S12, S28, and S29 mentioned considering software practitioners’ career aspirations in their recommendation systems as intended future improvements. Study S38 suggests including geographic regions, while Study S49 suggests demographic information. Hence, this reflects that there is an interest among researchers to explore human factors affecting their recommendation systems.
@@
    Transferring knowledge from one platform to another allows the new platform to benefit from pre-existing user preferences, task information, and interaction patterns [43] [44] [45]. For example, Studies S14, S15, and S27 mentioned addressing the “cold start” problem by transferring knowledge or merging more information from other platforms as future improvements to their systems. Researchers in other domains have researched transferring knowledge from other platforms for their recommendation systems. For example, Yan et al. investigated cross-platform social relationships and behavior information to tackle the cold-start problem in friend recommendation [46]. Specifically, they performed comprehensive data analysis to determine which information is most effectively transferable between platforms [46]. Pan et al. addressed the data sparsity issue in the recommendation systems by transferring knowledge about both users and items from auxiliary data sources [47]. Li et al. addressed the sparsity problem in individual rating matrices by learning a generative model from multiple related recommender systems [48].
@@
    Cross-platform recommendation systems provide users with personalized recommendations regardless of the platform they are signed up with. Only 4.6% of the primary studies were tested on multiple crowdsourcing platforms. When the recommendation systems are designed to cater to a specific platform, those systems may be constrained by the features and capabilities of the platform they are built for [53][54]. It may limit their flexibility and adaptability. Further platform-specific recommendation systems may contribute to user lock-in making it difficult for them to switch to alternatives.
    The flexibility of platform-independent solutions enables the recommendation system to cater to diverse user needs and preferences without being tied to a specific platform’s constraints. Interoperability in such solutions facilitates collaboration and data exchange between platforms, enabling more comprehensive and holistic recommendation strategies [55].
    However, developing cross-platform recommendation systems could be challenging as some required data may not be available on every platform in the same form [54] [56]. For example, the Topcoder platform offers money as a prize to the winners of their tasks whereas, GitHub does not provide such prizes for every task [31] [32]. Hence considering offered developer incentives as a recommendation factor in a cross-platform recommendation system is a complex task. On the other hand, when extending recommendation systems to work across multiple platforms, there’s a risk of sacrificing the unique features of each platform which could offer valuable insights for generating recommendations. For example, Studies S7 and S18 used the “GitHub’s forking a repository” feature to extract user preferences. When extending those recommendation systems to cater to multiple platforms, they may not use such unique platform features to extract data.
@@
    These further evaluations help to understand current gaps in the existing literature. Establishing standardized evaluation metrics, datasets, and benchmarking frameworks for comparing and assessing the performance of different recommendation approaches could facilitate reproducibility, transparency, and rigor in evaluating the effectiveness of software task recommendation systems across various settings and domains [57]. Further evaluations covering multiple aspects of recommendation systems, including technical metrics, real-world studies, and user feedback provide more comprehensive recommendations [58]. Evaluating recommendation models on larger and richer datasets helps validate their performance in real-world settings, ensuring that they can scale effectively and provide accurate recommendations across different contexts [59] [58].
@@
  • Integrating recommendation systems with developer workflow tools:

### v2

@@ 3 Results
The analysis and synthesis of the data extracted from our finalized 65 primary studies are reported in the following subsections. The results discussed in this section are derived by synthesizing the data with minimal interpretations directly extracted from the selected primary study set.
@@ 3.1 Demographic Data
Of the 65 primary studies selected for this SLR, 37 (56.9%) were conference papers, 26 (40%) were journal papers, and 2 (3.1%) were workshop papers. Figure 4 shows the yearly distribution and publication type distribution of the selected papers. The selected primary studies were published between 2011 and 2024 era even though we did not limit our search to a specific period. None of the included studies were published in 2012. There is an overall increase in the number of publications from 2011 to 2018. Then, in 2019 and 2020, there was a drop in the number of publications. The COVID-19 outbreak might have caused the sudden publication drop seen in 2020. However, it is not certain. Overall, 52% of the selected primary studies set were published from 2021 to 2024.
@@
Contributor’s social characteristics S8, S12, S34, S36, S38, S53, S65
Previous Contributions Skills, Expertise S2, S5, S10, S25, S27, S28, S29, S32, S37, S39, S40, S41, S42, S43, S45, S46, S52, S55, S64, S65
@@
Historical behavior sequences S5, S7, S8, S10, S17, S18, S19, S20, S21, S22, S23, S24, S32, S36, S47, S50, S52, S55, S60, S65
@@
Increasing performance, accuracy, and optimizing solutions: The performance, accuracy, and system optimizations in the selected primary studies were mainly achieved by using two or more characteristics mentioned in Section 3.2.1. For example, studies S8, S12, S34, S36, S38, S53, and S65 used software practitioner’s social connections with fellow software practitioners along with other characteristics to improve the accuracy of the systems. According to Study S12, the use of social metrics in their recommendation system improved the precision of the used models up to 15.82%. Further, some studies (e.g., S5, S7, S27) used hybrid recommendation systems (mentioned in section 3.2.2) to mitigate the limitations that occurred when using one recommendation approach.
@@
Study S28 solves the “cold start” problem by modelling worker’s preferences over multiple dimensions. Study S44 recommends tasks to newcomers by integrating repository history data (such as the number of prior commits made by the issue reporter at the time of the issue’s creation) and global GitHub profiles for all issue reporters and repository owners (e.g., stars received, commits). Study S59 recommends tasks to cold-start users by analyzing possible features and the characteristics of the newcomers’ chosen first issues. A new cross-platform recommendation approach for software crowdsourcing was introduced in S61 by transferring data and knowledge from other popular software crowdsourcing platforms. Study S61 adopts a transfer learning approach to solve the platform “cold start” problem.
@@
    Limitations in the approaches (7 papers): One of the key limitations identified in the proposed approaches is the lack of generalizability of the solution. The recommendation model in Study S3 primarily follows a content-based approach, limiting its ability to suggest issues to developers based solely on past contributions. It lacks the capability to provide more insightful recommendations that consider serendipity aspects across various components of the same product. Hence, the recommendations could be narrow. The NLP suggestions used in Study S12 might not work as expected in other languages like C++. Therefore, the approach cannot be utilized in projects using different programming languages. Whereas in Study S9, the projects within the scope of the study exhibit diverse practices, backgrounds, resources, and goals. These factors are not adequately considered during the model-building process. Hence, the recommendation system proposed in the study may not be applicable or valid across a broader range of projects with different contexts. Another key limitation is that some systems cannot recommend tasks to newcomers ( “cold start” problem). In Studies S2, S14, S15, S17 and S65, the proposed recommendation systems cannot recommend tasks to newcomers. As mentioned earlier in section 2.1, human factors play a major role in the software engineering context. However, only 24 primary studies have considered human factors in their recommendation systems. These human factors are analyzed in section 3.4. 39 primary studies ignored human factors in their recommendation systems. Some primary studies that ignored human factors are, S3, S6, S7, S10, S11, S14, S17, S20, 21, S22, etc.
@@ 3.3 RQ2: What are the common types of tasks that are recommended to software practitioners?
3.4 RQ3: To what extent are human factors used to recommend CSE tasks?
Human factors encompass personality traits, social and behavioral preferences, and sentiments, which are intrinsic to individuals and require different methods of measurement and analysis. To answer this research question, we analyzed the human factors (i.e., D09 and D10 in LABEL:table:7) currently used in the existing recommendation systems and the human factors suggested to be included as future improvements to the recommendation systems. Table 10 shows these human factors.
@@ 3.4 RQ3: To what extent are human factors used to recommend CSE tasks?
Studies S4, S57, and S58 investigated the impact of personality on task selection in crowdsourcing software development. In study S4, an empirical experiment was carried out to assess the impact of personality on task selection based on the three key components of a task: type, money, and time [24]. Study S57 was carried out to develop a task assignment algorithm for crowdsourcing software development that takes task types and personality types (e.g., Introversion and Extroversion, etc. ) into account. The proposed method functions in the same way as the open call format, but it incorporates personality-based categorization [25]. To identify the personality types of software practitioners, the Myers-Briggs Type Indicator (MBTI) was used [29]. MBTI is a popular tool used to analyze personality types in the workplace, which measures a person across four dimensions: energizing, attending, deciding, and living [30]. Study S58 was carried out to find the relationship between testing technique, tester, and personality type and identify which personality types of testers are more competent at a particular testing approach [31]. They also used the MBTI to identify the personality types [30].
@@
Demographic information, such as age, location, education, etc., are currently included in Studies S15 and S28, while Study S38 suggested considering the geographic region in the future. Study S15 modeled developer and project profiles by analyzing their basic information and mining user online behavior logs. Then they filtered out the projects based on the location, age, and education level requirements. This filtration process was done to speed up the recommendation process. It has been observed that workers from certain countries tend to prefer collaborating with task posters from specific other countries. For example, workers from India often prefer to work with task posters from the US and the UK [32]. Hence, Study S28 captured a worker’s preference for collaborating with task posters from certain countries. They extracted the worker’s location, education, etc., from the worker’s profile information.
@@
Skill preference S1, S5, S15, S16, S32, S59, S64 N/A
Social and behavioral preferences S8, S10, S43, S65 S5
@@ 3.5 RQ4: What platforms are currently being used to recommend crowdsourced software engineering tasks?
GitHub is the most popular platform used in the primary studies. It is used in 38 out of 65 studies (58.4% of the studies). It is a developer-based platform to build, scale, and deliver secure software [34]. Some of the Github projects used in the existing studies are JabRef, PowerToys, and Audacity. A further analysis of these projects is available in table LABEL:github_project_scales.
@@
RQ4 Summary: We observed that GitHub is the most popular 58.4% crowdsourced software engineering platform among the selected primary studies. Further, we observed that the majority (55%) of the GitHub projects that were used to recommend tasks were large-scale projects with more than 500 contributors. Further, Bugzilla, Topcoder, and Utest are some other popular CSE platforms (see Table LABEL:platforms_list).
@@ 3.6 RQ5: Which features of platforms can be used to recommend crowdsourced software engineering tasks?
Features of the platform include structured elements such as issue or task descriptions, pull requests, comments, labels, etc.; these are tangible, system-generated artifacts within the development platform. Table 13 shows the features and the different attributes of the features of crowdsourcing platforms used to extract data.
@@
User activities such as starring, forking, watching, and bidding were extracted in 24 primary studies. This is the second most popular feature used in primary studies. For example, Study S7 extracted starring and forking repositories in the GitHub platform to identify user preferences.
Pull requests are the third most popular feature used to extract different attributes. It was used in 18 primary studies. Commit message text, names of the files changed in the pull request, and pull request comments were extracted from the pull requests. For example, Study S1 extracted the name of the files changed, and commit messages from the pull requests to identify the relevant domain for the issues.
Comments were used in 14 primary studies. commenter’s name and the text were extracted from the comments. Study S12 extracted commenter names to get the total number of commenters which they used as a social metric in their recommendation system.
@@
RQ5 Summary: Issue or task description is the most popular feature used to extract data in primary studies. Further, pull requests, comments, labels, tags, etc., are other popular features used to extract data (see Table 13). From those features, different attributes, such as commit message, task type, APIs used, etc, were extracted.
@@
User activities Fork, star, watch, bid, etc S5, S7, S16, S18, S19, S20, S21, S22, S23, S24, S32, S34, S35, S36, S44, S47, S50, S51, S56, S57, S62, S63, S64,S65%
Pull Requests Commit messages S1, S10, S11, S12, S18, S20, S33, S36, S41, S44, S53, S59, S65
@@ 4 Threats to Validity
Internal validity. To mitigate threats to internal validity, we adhered to a predefined SLR protocol. The search string was iteratively refined and tested before execution in order to optimize the results. Further, we developed our search string according to PICOC criteria, which enhances its comprehensiveness. To minimiz bias in the study selection process, we followed a multi-step filtering process. The first two authors validated the inclusion and exclusion criteria on a small subset of primary studies before applying them broadly. Papers were screened in several rounds: initially based on titles, abstracts, and keywords followed by a brief reading, and finally, a detailed review. Any disagreements regarding inclusion were resolved through discussions between the first and second authors. For data extraction, we developed a structured data extraction form (see Table 7) to ensure consistency. Since the first author conducted most of the extraction, regular discussions were held with the second author to clarify uncertainties and resolve discrepancies. A subset of the extracted data was further verified by the second and third authors.
Construct validity. To reduce threats to construct validity, we searched across six major digital libraries and employed two complementary search strategies: automatic database searches and snowballing. The selected primary studies provide highly relevant solutions to recommend a CSE task to a software practitioner. To further refine our selection process, we conducted several rounds of discussions to improve the inclusion and exclusion criteria.
Conclusion validity. To minimize biases in data synthesis and analysis, we adopted a collaborative approach. All authors participated in multiple rounds of discussions to determine the most effective way to categorize and present findings. This iterative process helped reduce subjective interpretations and ensured that our conclusions were well-grounded in the extracted data.
External validity. External validity reflects the generalizability of our findings. A potential limitation is the scope of the literature included. Our study relied on papers retrieved from six major digital libraries, including Scopus, which is known for its extensive indexing capabilities. However, studies outside these databases may not have been captured. To mitigate this, we employed snowballing techniques, manually reviewing references from selected papers to identify additional relevant studies. Further, we did not impose any time limitation on our search to capture all developments in the area. To improve the generalizability of our findings, future work could incorporate additional sources.
5 Related Work
Zhen et al. [12] conducted an SLR that focuses on the software tasks in the crowdsourcing domain at a high level, along with the tasks in non-software engineering domains such as data entry, writing, editing, etc. The study examined the impact of task assignment in crowdsourcing, including the consequences of assigning a task to an appropriate worker and an inappropriate worker. The study highlights that improper task-worker matching not only affects result quality but also wastes valuable resources such as time, money, and client trust, underscoring the need for an optimal task assignment model. It emphasized the application of crowdsourcing in software engineering and analyzed existing methods for effective task allocation. Further, it identified 52 existing crowdsourcing platforms such as Amazon MechanicalTurk (AMT)444 https://www.mturk.com/, Topcoder555 https://www.topcoder.com, CastingWords666 http://castingwords.com/. Our SLR provides a comprehensive, in-depth analysis of personalized task recommendations to software practitioners in crowdsourcing platforms, focusing on recommending a task to a particular software practitioner in a crowdsourcing platform. Further, the existing SLR is conducted on the studies published from 2010 to 2019, whereas our SLR has no time duration restriction. Our search string significantly differs from the one used in [12]. They identified 120 relevant studies and selected 70 most relevant studies for their SLR. Only one primary study overlaps with the selected primary studies of our SLR. Regarding the research questions, our RQ1, RQ2, and RQ4 partially overlap with the existing literature review’s RQ3, RQ2, and RQ4, respectively. Moreover, we focus on assigning a suitable task to a particular software practitioner, whereas the majority of primary studies in the existing SLR focus on assigning a practitioner to a task. Since the existing SLR is a high-level study on the software engineering domain, the results include those unrelated to the software engineering field. Hence, there is a significant difference between our SLR and the existing SLR.
Mao et al. [1] performed a survey that focuses on how crowdsourcing is used in different stages of the software development life cycle along with its taxonomy. Further, it provides definitions for crowdsourced software engineering, crowdsourcing practices in software engineering, commercial platforms, and relevant case studies. The study highlighted that while many previous studies have cited benefits such as low cost and quick delivery, there is a lack of comparative research that validates these advantages against traditional development methods. It also highlighted the significance of quality control in crowdsourced projects. Further, the findings of the study emphasize the challenges of managing communication, expectations, and workflows among crowd contributors with varying skill levels and motivations. It also highlights concerns about intellectual property rights and the need for clear guidelines to protect both requesters and contributors. Additionally, the study notes that cultural and regional differences can affect collaboration, making it crucial to understand these differences for effective global crowd management and alignment with project goals. However, the survey is not based on any research questions as it is a broad analysis. They conducted the survey on the studies published from January 2006 to March 2015, whereas our SLR has no time duration restriction. They surveyed 210 publications out of 476 studies identified during the initial stage. There are no overlapping primary studies with our selected primary studies set for this SLR. Our search string significantly differs from the search string used in this survey. RQ4 in our SLR overlaps with the section “4.1 commercial platforms” in the survey. However, our SLR is focused on recommending crowdsourced software engineering tasks to software practitioners, whereas this survey is a broad analysis of crowdsourced software engineering. Hence, there is a significant difference in this survey with our SLR.
@@ 6 Discussion and Future Research Directions
    Software is developed by humans, read by humans, and maintained by humans [41]. Human factors play a crucial role in software development [42]. Human factors affect software development team members, their activities, and ultimately, the quality of the product [15]. Ockiya and Lock [43] discovered a connection between human factors in remote software teams and their success or failure. Gaining a deeper understanding of this relationship and managing it effectively could help alleviate the challenges faced by remote teams. Some major challenges in CSE include ensuring effective communication and coordination, handling knowledge sharing, fostering motivation, building a network of volunteers, establishing trust, and facilitating team growth [4, 44, 45].
    Human factors have been integrated into some existing primary studies, and several studies plan to incorporate them as future enhancements to their recommendation systems. We analyzed the human factors currently being used or suggested to be used in the existing systems in Table 10. For example, Studies S9 and S59 used software practitioners’ sentiment data in their recommendation system. Study S29 considers software practitioners’ motivation when recommending tasks, whereas Studies S18 and S28 plan to consider motivation as a future improvement. Studies S12, S28, and S29 mentioned considering software practitioners’ career aspirations in their recommendation systems as intended future improvements. Study S38 suggests including geographic regions, while Study S49 suggests demographic information. Hence, this reflects that there is an interest among researchers to explore human factors affecting their recommendation systems.
@@
    As new software crowdsourcing platforms often lack sufficient historical data on developer behaviour, existing recommendation algorithms struggle to effectively match developers with new tasks [49]. Among the existing algorithms, collaborative filtering (CF) has gained significant recognition due to its versatility, interpretability, and independence from content-based features. However, it is also highly susceptible to data sparsity and the cold start problem. To address these challenges, CF algorithms with transfer learning have received considerable attention [50, 51]. By leveraging knowledge from data-rich auxiliary domains and transferring it to target domains with limited data, these systems help mitigate data sparsity issues and enhance recommendation accuracy. This transfer of knowledge enables new platforms to leverage existing user preferences, task information, and interaction patterns, thus enhancing their overall performance and user experience [52] [53] [54]. Given that contributors in CSE come from diverse backgrounds with varying levels of experience, leveraging prior data from different platforms can help match developers to projects more effectively.
    For example, Studies S14, S15, and S27 mentioned addressing the “cold start” problem by transferring knowledge or merging more information from other platforms as future improvements to their systems. Researchers in other domains have researched transferring knowledge from other platforms for their recommendation systems. For example, Yan et al. investigated cross-platform social relationships and behavior information to tackle the cold-start problem in friend recommendation [55]. Specifically, they performed comprehensive data analysis to determine which information is most effectively transferable between platforms [55]. Pan et al. addressed the data sparsity issue in the recommendation systems by transferring knowledge about both users and items from auxiliary data sources [56]. Li et al. addressed the sparsity problem in individual rating matrices by learning a generative model from multiple related recommender systems [57].
@@
    Cross-platform recommendation systems provide users with personalized recommendations regardless of the platform they are signed up with.These systems enable a broader reach and cater to the diverse needs of users who participate in crowdsourcing platforms. Only 4.6% of the primary studies were tested on multiple crowdsourcing platforms. When the recommendation systems are designed to cater to a specific platform, those systems may be constrained by the features and capabilities of the platform they are built for [62][63]. It may limit their flexibility and adaptability. Further platform-specific recommendation systems may contribute to user lock-in where users may feel constrained to stay on a single platform due to a lack of integration with other services or options. For example, Xu et al. [64] proposed a cross-platform developer recommendation algorithm based on feature matching. Experimental results demonstrate that the proposed algorithm outperforms existing recommendation algorithms across various evaluation metrics.
    The flexibility of platform-independent solutions enables the recommendation system to cater to diverse user needs and preferences without being tied to a specific platform’s constraints. The interoperability of these solutions facilitates collaboration and data exchange between platforms, enabling more comprehensive and holistic recommendation strategies [65]. This is particularly important in crowdsourcing, where users often contribute to projects across different platforms and bring a variety of skills, motivations, and preferences to the table.
    However, developing cross-platform recommendation systems could be challenging as some required data may not be available on every platform in the same form [63] [66]. For example, the Topcoder platform offers money as a prize to the winners of their tasks, whereas GitHub does not provide such prizes for every task [34] [35]. Hence, considering offered developer incentives as a recommendation factor in a cross-platform recommendation system is a complex task. On the other hand, when extending recommendation systems to work across multiple platforms, there’s a risk of sacrificing the unique features of each platform, which could offer valuable insights for generating recommendations. For example, Studies S7 and S18 used the “GitHub’s forking a repository” feature to extract user preferences. When extending those recommendation systems to cater to multiple platforms, they may not use such unique platform features to extract data.
@@
    These further evaluations help to understand current gaps in the existing literature. Establishing standardized evaluation metrics, datasets, and benchmarking frameworks for comparing and assessing the performance of different recommendation approaches could facilitate reproducibility, transparency, and rigor in evaluating the effectiveness of software task recommendation systems across various settings and domains [67]. Further evaluations covering multiple aspects of recommendation systems, including technical metrics, real-world studies, and user feedback, provide more comprehensive recommendations [68]. Evaluating recommendation models on larger and richer datasets helps validate their performance in real-world settings, ensuring that they can scale effectively and provide accurate recommendations across different contexts [69] [68].
@@
  • Inegrating recommendation systems with development workflow tools: